# Geodesic Distance Between Graphs: A Spectral Metric for Assessing the Stability of Graph Neural Networks

### v1

@@ 7.4 Partial Node Features
Table 4: Comparison of Correlation with GNN outputs and distance metrics with Partial Node Features
Figure 4: Percentage of the original GGD using numbers of extreme Eigenvalues
@@ 7.5 GGD Approximation using Extreme Eigenvalues
Table 5: Performance of GGD using Extreme Eigenvalues only
@@
The largest or smallest eigenvalues correspond to the most dominant mismatches in graph cuts and effective resistance distances, contributing most to the total GGD value. Similarly, the second largest and smallest eigenvalues correspond to the next significant mismatched cuts. In our experiment, we obtain approximate GGDs using a few extreme eigenvalue pairs and compare them with the ground truth. Figure 4 illustrates the relative accuracy of the approximate GGDs, in which we observe that the top four pairs of extreme eigenvalues contribute $80\%$ of the total GGD values. In addition, we conduct the SVC classification task and GNN correlation study using GGD with only $2$ and $4$ extreme eigenvalue pairs, respectively and present the associated findings in Table 5.
@@ Appendix A Appendix
A.1 Detailed Proofs Showing GGD is a Pseudometric
A.1.1 Identity Property
@@
$GGD(G,G)=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(\mathcal{L}^{-1}\mathcal{L}))\right]^{1/2}=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(I))\right]^{1/2}$
The identity matrix has only one eigenvalue, which is 1. So, $GGD(G,G)=\left[\log^{2}(1)\right]^{1/2}=0$
A.1.2 Positivity Property
@@
$GGD(G_{1},G_{2})=\left[\log^{2}(\lambda_{1})+\log^{2}(\lambda_{2})+\log^{2}(\lambda_{3})+\ldots+\log^{2}(\lambda_{n})\right]^{1/2}$
@@ A.1.3 Symmetry Property
$GGD(G_{1},G_{2})=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i})\right]^{1/2}$
@@
$GGD(G_{2},G_{1})=\left[\sum_{i=1}^{n}\log^{2}(\frac{1}{\lambda_{i}})\right]^{1/2}$
Now, $\log\left(\frac{1}{\lambda_{i}}\right)=-\log(\lambda_{i})$; so, $\log^{2}\left(\frac{1}{\lambda_{i}}\right)=\log^{2}(\lambda_{i})$
So, we can conclude $GGD(G_{1},G_{2})=GGD(G_{2},G_{1})$
A.1.4 Triangle Inequality
@@
Now, The Frobenius norm $\|X\|_{F}$ is the geodesic length at $d(\exp X,I)=\|X\|_{F}$ Bonnabel and Sepulchre [2010]. Hence at identity, $d(\mathcal{L},I)=\|\log\mathcal{L}\|_{F}$.
From Bonnabel and Sepulchre [2010], You and Park [2021] we get,
$GGD(G_{1},G_{2})=GGD\left(G_{1}^{-1/2}G_{2}G_{1}^{-1/2},I\right)=\left\|\log\left(\mathcal{L}_{1}^{-1/2}\mathcal{L}_{2}\mathcal{L}_{1}^{-1/2}\right)\right\|_{F}=\left\|\log\left(\mathcal{L}_{1}^{-1}\mathcal{L}_{2}\right)\right\|_{F}$ (16)
@@
$\mathcal{L}_{1}^{-1}\mathcal{L}_{3}=\mathcal{L}_{1}^{-1}(\mathcal{L}_{2}\mathcal{L}_{2}^{-1})\mathcal{L}_{3}=(\mathcal{L}_{1}^{-1}\mathcal{L}_{2})(\mathcal{L}_{2}^{-1}\mathcal{L}_{3})$
@@
$\|\mathcal{L}_{1}^{-1}\mathcal{L}_{3}\|=\|(\mathcal{L}_{1}^{-1}\mathcal{L}_{2})(\mathcal{L}_{2}^{-1}\mathcal{L}_{3})\|\leq\|\mathcal{L}_{1}^{-1}\mathcal{L}_{2}\|\|\mathcal{L}_{2}^{-1}\mathcal{L}_{3}\|$
@@
$\|\log(\mathcal{L}_{1}^{-1}\mathcal{L}_{3})\|\leq\|\log(\mathcal{L}_{1}^{-1}\mathcal{L}_{2})\|+\|\log(\mathcal{L}_{2}^{-1}\mathcal{L}_{3})\|$
@@
$GGD(G_{1},G_{3})\leq GGD(G_{1},G_{2})+GGD(G_{2},G_{3})$
A.2 Graph Matching Recovery
Given symmetric matrices $A$, $B$ and $Z$ from the Gaussian Wigner model, where $B_{\pi^{*}}=A+\sigma Z$, there exist constants $c,c^{\prime}>0$ such that if $1/n^{0.1}\leq\eta\leq c/\log n$ and $\sigma\leq c^{\prime}\eta$, then with probability at least $1-n^{-4}$, GRAMPA Algorithm correctly recovers the permutation matrix $\pi^{*}$.
From 6, the similarity matrix $\widehat{X}$ is defined as:
$\widehat{X}=\sum_{i,j=1}^{n}\frac{u_{i}u_{i}^{T}\mathbf{J}v_{j}v_{j}^{T}}{(\zeta_{i}-\mu_{j})^{2}+\eta^{2}},$
where $u_{i}$ and $v_{j}$ are eigenvectors of $A$ and $B$ respectively, and $\zeta_{i}$ and $\mu_{j}$ are their corresponding eigenvalues. $\mathbf{J}$ is a all-one matrix, and $\eta$ is the tuning parameter.
Now, the proof is divided into two parts:
@@ A.2 Graph Matching Recovery
In a noiseless situation, means replacing B with A, similarity matrix $\widehat{X}^{*}$ is defined as:
$\widehat{X}^{*}=\sum_{i,j=1}^{n}\frac{u_{i}u_{i}^{T}\mathbf{J}u_{j}u_{j}^{T}}{(\zeta_{i}-\zeta_{j})^{2}+\eta^{2}}$ (17)
For some constants $C,c>0$, if $1/n^{0.1}<\eta<c/\log n$, then with probability at least $1-5n^{-5}$ for large $n$, it can be proved that the diagonal components of $\widehat{X}^{*}$ are dominant by showing Fan et al. [2020]:
$\min_{i\in[n]}(\widehat{X}^{*})_{ii}>\frac{1}{3\eta^{2}},\quad\text{and}\quad\max_{i,j\in[n]:i\neq j}(\widehat{X}^{*})_{ij}<C\left(\frac{\sqrt{\log n}}{\eta^{3/2}}+\frac{\log n}{\eta}\right)$ (18)
@@
The difference between the similarity matrix $X$ in the presence of noise and the noiseless situation is bounded. If $\eta>1/n^{0.1}$, then for a constant $C>0$, with probability at least $1-2n^{-5}$ for large $n$, it can be shown:
$\max_{i,j\in[n]}|\widehat{X}_{ij}-(\widehat{X}^{*})_{ij}|<C\sigma\left(\frac{1}{\eta^{3}}+\frac{\log n}{\eta^{2}}\left(1+\frac{\sigma}{\eta}\right)\right)$ (19)
Assuming Lemma A.1 and A.2, for some $c,c^{\prime}>0$ sufficiently small, and by setting $\eta<c/\log n$ and $\sigma<c^{\prime}\eta$, the algorithm ensures that all diagonal entries of $\widehat{X}$ are larger than all off-diagonal entries, thereby achieving exact recovery.
A.3 Effect of tuning parameter $\eta$ on Graph Matching
@@
In the original work, it is suggested that the regularization parameter $\eta$ needs to be chosen so that $\sigma\vee n^{-0.1}\lesssim\eta\lesssim 1/\log n$ Fan et al. [2020]. It is also mentioned that for practical cases, computing permutation matrix for different values of $\eta$ in an iterative way can result in better accuracy. The GRAMPA uses $\eta=0.2$ for all their experiments Fan et al. [2020].
@@
We used a few values of $\eta$ in the classification problem using MUTAG dataset and got that the best accuracy is obtained at $\eta=0.5$. In Figure 5, the performance of the tuning parameter is demonstrated.
Figure 5: Classification Accuracy vs GRAMPA tuning parameter
A.4 Experimental Setup

### v2

@@ 7.4 Partial Node Features
Table 4: Comparison of correlation with GNN outputs and distance metrics with partial node features.
@@
Figure 4: Percentage of the original GGD using numbers of extreme eigenvalues.
@@ 7.5 GGD Approximation using Extreme Eigenvalues
Table 5: Performance of GGD using extreme eigenvalues only.
@@
The largest and the smallest eigenvalues correspond to the most dominant mismatches in graph cuts and effective resistance distances, contributing the most to the total GGD value. Similarly, the second largest and smallest eigenvalues correspond to the next significant mismatched cuts. In our experiment, we obtain approximate GGDs using a few extreme eigenvalue pairs and compare them with the ground truth. Figure 4 illustrates the relative accuracy of the approximate GGDs, in which we observe that the top four pairs of extreme eigenvalues contribute $80\%$ of the total GGD values. In addition, we conduct the SVC classification task and GNN correlation study using GGD with only $2$ and $4$ extreme eigenvalue pairs, respectively, and present the associated findings in Table 5.
@@ Appendix A Appendix
A.1 Graph Adjacency and Laplacian Matrices
For an undirected graph $G=(V,E,w)$, where $V$ represents the set of nodes (vertices), $E$ represents the set of edges, and $w$ denotes the associated edge weights, the adjacency matrix $A$ is defined as follows:
$A(i,j)=\begin{cases}w(i,j),&\text{ if }(i,j)\in E.\\ 0,&\text{ otherwise. }\end{cases}$ (16)
Let $D$ denote the diagonal matrix where $D(i,i)$ is equal to the (weighted) degree of node $i$. The graph Laplacian matrix is then given by $L=D-A$. The rank of the Laplacian matrix of a graph $G$ is $n-c(G)$, where $n$ is the number of nodes and $c(G)$ is the number of connected components in the graph. For a connected graph, this implies that the rank of the Laplacian matrix is $n-1$, meaning Laplacian matrices are not full-rank (Bondy et al., 1976).
A.2 Riemannain Manifold
A manifold is a type of topological space that resembles Euclidean space in small, local regions around each point. In other words, for every point on a manifold, there is a neighborhood that is similar to a flat multidimensional space. A Riemannian manifold is a type of manifold equipped with a smoothly varying inner product on the tangent spaces at each point. This means that for every point on the manifold, the tangent space has a way of measuring distances and angles, and these measurements change smoothly from point to point (Lee, 2018). In simpler terms, a Riemannian manifold is a smooth, curved space that locally behaves like Euclidean space but has its own geometric properties, such as how distances, angles, and volumes are defined. These properties are determined by a Riemannian metric, which generalizes the concept of measuring lengths and angles in flat space to curved spaces (Lee, 2018).
A Riemannian manifold can have curvature, unlike a flat space. This curvature allows the study of geometric shapes ranging from spheres and cylinders to more abstract surfaces. The Riemannian structure enables us to compute geodesics, volumes, and various types of curvature. This makes Riemannian manifolds fundamental in fields like differential geometry and physics, and increasingly important in data science, where curved spaces are used to model complex datasets (You & Park, 2021).
A.3 Effective Resistance in Graph Theory
Effective resistance, also known as resistance distance, is a concept in graph theory that measures the electrical resistance between two nodes in a weighted or unweighted graph represented as a network of resistors (Ghosh et al., 2008). It draws an analogy between electrical networks and graphs, helping to quantify how easily current can flow between two nodes, where the edges are treated as resistors. The effective resistance between nodes provides insight into the connectivity between the network. This means two nodes with lower effective resistance values have higher connectivity (Ellens et al., 2011).
A.4 Detailed Proofs Showing GGD is a Metric
A.4.1 Identity Property
@@
$GGD(G,G)=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(\mathcal{L}^{-1}\mathcal{L}))\right]^{1/2}=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(I))\right]^{1/2}.$
The identity matrix has only one eigenvalue, which is 1. So, $GGD(G,G)=\left[\log^{2}(1)\right]^{1/2}=0$. ∎
A.4.2 Positivity Property
@@
$GGD(G_{1},G_{2})=\left[\log^{2}(\lambda_{1})+\log^{2}(\lambda_{2})+\log^{2}(\lambda_{3})+\ldots+\log^{2}(\lambda_{n})\right]^{1/2}.$
@@ A.4.3 Symmetry Property
$GGD(G_{1},G_{2})=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i})\right]^{1/2}.$
@@
$GGD(G_{2},G_{1})=\left[\sum_{i=1}^{n}\log^{2}(\frac{1}{\lambda_{i}})\right]^{1/2}.$
Now, $\log\left(\frac{1}{\lambda_{i}}\right)=-\log(\lambda_{i})$; so, $\log^{2}\left(\frac{1}{\lambda_{i}}\right)=\log^{2}(\lambda_{i})$.
So, we can conclude $GGD(G_{1},G_{2})=GGD(G_{2},G_{1})$. ∎
A.4.4 Triangle Inequality
@@
Now, The Frobenius norm $\|X\|_{F}$ is the geodesic length at $d(\exp X,I)=\|X\|_{F}$ (Bonnabel & Sepulchre, 2010). Hence at identity, $d(\mathcal{L},I)=\|\log\mathcal{L}\|_{F}$.
From (Bonnabel & Sepulchre, 2010; You & Park, 2021) we get,
$GGD(G_{1},G_{2})=GGD\left(G_{1}^{-1/2}G_{2}G_{1}^{-1/2},I\right)=\left\|\log\left(\mathcal{L}_{1}^{-1/2}\mathcal{L}_{2}\mathcal{L}_{1}^{-1/2}\right)\right\|_{F}=\left\|\log\left(\mathcal{L}_{1}^{-1}\mathcal{L}_{2}\right)\right\|_{F}.$ (17)
@@
$\mathcal{L}_{1}^{-1}\mathcal{L}_{3}=\mathcal{L}_{1}^{-1}(\mathcal{L}_{2}\mathcal{L}_{2}^{-1})\mathcal{L}_{3}=(\mathcal{L}_{1}^{-1}\mathcal{L}_{2})(\mathcal{L}_{2}^{-1}\mathcal{L}_{3}).$
@@
$\|\mathcal{L}_{1}^{-1}\mathcal{L}_{3}\|=\|(\mathcal{L}_{1}^{-1}\mathcal{L}_{2})(\mathcal{L}_{2}^{-1}\mathcal{L}_{3})\|\leq\|\mathcal{L}_{1}^{-1}\mathcal{L}_{2}\|\|\mathcal{L}_{2}^{-1}\mathcal{L}_{3}\|.$
@@
$\|\log(\mathcal{L}_{1}^{-1}\mathcal{L}_{3})\|\leq\|\log(\mathcal{L}_{1}^{-1}\mathcal{L}_{2})\|+\|\log(\mathcal{L}_{2}^{-1}\mathcal{L}_{3})\|.$
@@
$GGD(G_{1},G_{3})\leq GGD(G_{1},G_{2})+GGD(G_{2},G_{3}).$
A.5 Graph Matching Recovery
Given symmetric matrices $A_{1}$, $A_{2}$ and $Z$ from the Gaussian Wigner model, where $A_{2\pi^{*}}=A_{1}+\sigma Z$, there exist constants $c,c^{\prime}>0$ such that if $1/n^{0.1}\leq\eta\leq c/\log n$ and $\sigma\leq c^{\prime}\eta$, then with probability at least $1-n^{-4}$ for all large n, the matrix $\widehat{X}$ in equation 6 satisfies,
$\min_{i\in[n]}\widehat{X}_{i,\pi^{*}(i)}>\max_{i,j\in[n]:j\neq\pi^{*}(i)}\widehat{X}_{ij}$ (18)
and hence, the GRAMPA algorithm correctly recovers the permutation estimation matrix $\pi^{*}$.
Now, this proof is divided into two parts:
@@ A.5 Graph Matching Recovery
In a noiseless situation, means replacing $A_{2}$ with $A_{1}$, similarity matrix $\widehat{X}^{*}$ is defined as:
$\widehat{X}^{*}=\widehat{X}(A_{1},A_{1})=\sum_{i,j=1}^{n}\frac{u_{i}u_{i}^{T}\mathbf{J}u_{j}u_{j}^{T}}{(\zeta_{i}-\zeta_{j})^{2}+\eta^{2}}.$ (19)
For some constants $C,c>0$, if $1/n^{0.1}<\eta<c/\log n$, then with probability at least $1-5n^{-5}$ for large $n$, it can be proved that the diagonal components of $\widehat{X}^{*}$ are dominant by showing (Fan et al., 2019):
$\min_{i\in[n]}(\widehat{X}^{*})_{ii}>\frac{1}{3\eta^{2}}$ (20)
and
$\max_{i,j\in[n]:i\neq j}(\widehat{X}^{*})_{ij}<C\left(\frac{\sqrt{\log n}}{\eta^{3/2}}+\frac{\log n}{\eta}\right).$ (21)
@@
The difference between the similarity matrix $X$ in the presence of noise and the noiseless situation is bounded. If $\eta>1/n^{0.1}$, then for a constant $C>0$, with probability at least $1-2n^{-5}$ for large $n$, it can be shown (Fan et al., 2019):
$\max_{i,j\in[n]}|\widehat{X}_{ij}-(\widehat{X}^{*})_{ij}|<C\sigma\left(\frac{1}{\eta^{3}}+\frac{\log n}{\eta^{2}}\left(1+\frac{\sigma}{\eta}\right)\right).$ (22)
Assuming lemma A.1 and A.2, for some $c,c^{\prime}>0$ sufficiently small, and by setting $\eta<c/\log n$ and $\sigma<c^{\prime}\eta$, the algorithm ensures that the right sides of both equations 21 and 22 are at most $1/(12\eta^{2})$. Then when $\pi^{*}=id$ (the identity permutation), these lemmas combine to imply:
$\min_{i\in[n]}\widehat{X}_{ii}>\frac{1}{4\eta^{2}}>\frac{1}{6\eta^{2}}>\max_{i,j\in[n]:i\neq j}\widehat{X}_{ij}$ (23)
with probability at least $1-n^{-4}$. So, all diagonal entries of $\widehat{X}$ are larger than all off-diagonal entries, thereby achieving exact recovery (Fan et al., 2019).
A.6 Choosing $\epsilon$ for converting Laplacians to SPD matrices
Laplacian matrices are symmetric positive semi-definite (SPSD) matrices. To convert these to symmetric positive definite (SPD) matrices, we added a diagonal matrix with very small values ($\epsilon$). We used $0.0001$ as the small value ($\epsilon$) in our experiments. When working with Laplacian matrices of a weighted or unweighted graph, values significantly smaller than the edge weights of that graph have a minimal effect on the transformation. We conducted additional experiments with different small values and included the results in Tables 6 and 7. In Table 6, we observed that in our specific case with the MUTAG graph dataset (Morris et al., 2020), where all graphs are unweighted, any value less than 0.001 has an almost negligible influence on the performance of the graph classification task. Additionally, when using values equal to or less than 0.001, the GGD value remains almost the same, as shown in Table 7.
Table 6: Classification accuracy using MUTAG dataset with different values of $\epsilon$.
Value of $\epsilon$ Classification accuracy Value of $\epsilon$ Classification accuracy
0.1 76.38 $\pm$ 7.9 1e-4 85.96 $\pm$ 5.26
5e-2 79 $\pm$ 6.8 1e-5 84.21 $\pm$ 5.26
1e-2 79 $\pm$ 5.26 1e-6 85.96 $\pm$ 5.26
5e-3 81.57 $\pm$ 7.9 1e-7 85.96 $\pm$ 7.9
1e-3 81.57 $\pm$ 7.9
Table 7: GGD values using MUTAG dataset for different values of $\epsilon$.
Value of Normalized GGD of a random Average normalized
$\epsilon$ graph pair (MUTAG[85], MUTAG[103]) GGD of 1000 pairs
0.1 0.712 0.727
5e-2 0.827 0.834
1e-2 0.952 0.959
5e-3 0.978 0.979
1e-3 0.996 0.995
1e-4 0.9996 0.9995
1e-5 0.99995 0.99996
1e-6 0.999996 0.999996
1e-7 1 1
A.7 Comparison of Two Different Riemannian Metrics for SPD Matrices
The two most commonly used Riemannian metrics on the SPD manifold are the Affine Invariant Riemannian Metric (AIRM) and the Log-Euclidean Riemannian Metric (LERM) (Ilea et al., 2018; Thanwerdas & Pennec, 2023; Chen et al., 2024). AIRM is a Riemannian metric that remains invariant under affine transformations, meaning the metric is unaffected when matrices are transformed by any invertible operation. The geodesic distance between two SPD matrices, A and B, using AIRM is given by (You & Park, 2021; Moakher, 2005):
$d_{\text{AIRM}}(A,B)=\|\log(A^{-1}B)\|_{F}=\left[\sum_{i=1}^{n}\log^{2}(\lambda_{i}(A^{-1}B))\right]^{1/2}.$ (24)
On the other hand, LERM addresses some of the computational complexity challenges associated with AIRM by mapping SPD matrices to an Euclidean space through the matrix logarithm operation. In this Euclidean space, computations are simplified. The geodesic distance between two SPD matrices, A and B, using LERM is defined as (Huang et al., 2014):
$d_{\text{LERM}}(A,B)=\|\log(A)-\log(B)\|_{F}.$ (25)
In this paper, we primarily used AIRM to compute geodesics because of its stronger theoretical foundation and its ability to better explain graph cut mismatches. However, for comparison, we also conducted experiments using LERM. Figure 5 shows that the Graph Geodesic Distances computed using the LERM metric are highly correlated with those obtained using AIRM, though the GGD using AIRM demonstrates better performance overall. A detailed performance comparison of these two metrics is provided in Table 8.
Table 8: Comparison between Riemannian metrics for GGD calculation.
Riemannian metric Correlation with GNN output Classification accuracy
Affine-Invariant 0.7786 86±7.5%
Log-Euclidean 0.7634 84.38%
Figure 5: GGD between graph pairs using AI and LE Riemannian metric.
A.8 Using Normalized Laplacians for GGD Calculation
In many existing studies, the normalized Laplacian matrix is commonly used to study spectral graph properties (Chung, 1997). The normalized Laplacian matrix of a graph $G$ is defined as: $L_{norm}=I-A_{norm}$, where $A_{norm}$ is the normalized adjacency matrix. The normalized adjacency matrix is expressed as (Chung, 1997):
$A_{norm}=D^{-1/2}AD^{-1/2},$ (26)
where $D$ represents the diagonal degree matrix, and $A$ denotes the adjacency matrix of the graph.
Form equation 26, we can derive:
$L_{norm}=I-D^{-1/2}AD^{-1/2}=D^{-1/2}(D-A)D^{-1/2}=D^{-1/2}LD^{-1/2}.$ (27)
Similar to the Laplacian matrices of graphs, normalized graph Laplacian matrices are also symmetric and positive semi-definite. Therefore, it is necessary to add small values to the diagonal elements of these matrices. However, our experiments reveal that the GGD calculation is highly sensitive to this small value ($\epsilon$), resulting in significant fluctuations across different values, as demonstrated in Table 9. Additionally, the geodesic distances computed with the modified normalized Laplacian matrices exhibit poor accuracy in both classification tasks and their correlation with GNN outputs.
Table 9: Effect of epsilon in the calculation of GGD using normalized Laplacian matrices.
Value of $\epsilon$ 0.01 0.001 0.0001 0.00001
GGD using Laplacian matrices 16.235 16.775 16.832 16.838
GGD using normalized Laplacian matrices 384.1 254.44 188.345 165.332
A.9 Effect of tuning parameter $\eta$ on Graph Matching
@@
In the original work, it was suggested that the regularization parameter $\eta$ needs to be chosen so that $\sigma\vee n^{-0.1}\lesssim\eta\lesssim 1/\log n$ (Fan et al., 2020). It is also mentioned that for practical cases, computing permutation matrix for different values of $\eta$ in an iterative way can result in better accuracy. The GRAMPA uses $\eta=0.2$ for all their experiments (Fan et al., 2020).
@@
We used a few values of $\eta$ in the classification problem using the MUTAG dataset and got that the best accuracy is obtained at $\eta=0.5$. In Figure 6, the performance of the tuning parameter is demonstrated.
Figure 6: Classification accuracy vs GRAMPA tuning parameter.
A.10 Experimental Setup